OpenStudy (anonymous):

given y=tanxy+x^2, find dy/dx

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

y' = sec^2 (xy)* [ y + xdy/dx) + 2x

OpenStudy (anonymous):

thank you laplace2!

OpenStudy (anonymous):

y' = sec^2 ( xy)*y + sec^2 (xY) * x *y' + 2x , distributing

OpenStudy (anonymous):

collect like terms

OpenStudy (anonymous):

y ' [ 1 - sec^(xy) * x ] = sec^2 (xy)*y + 2x

OpenStudy (anonymous):

y' = ( sec^2 (xy)*y + 2x ) / ( 1 - sec^2(xy)*x)